A metal wire of resistivity 64x 10-6 ohm cm and length 198cm has resistance of 7 ohm. Calculate its radius.
1
Expert's answer
2012-07-12T08:28:43-0400
R = a * V = a * l / (Pi * r^2)
r = sqrt((a * l / R) / Pi) = sqrt((64x10^-6 *
198 / 7) / 3.14159) = 0.024005 cm